OpenStudy (anonymous):

1-cosx=-sinx x∈[-180, 180] I moved the -sinx over to the other side so now I have -cosx+sinx+1=0 but I'm not sure where to go from there. The answers are 0 and -90 degrees but I don't know how to get there. Help is much appreciated, thanks!

OpenStudy (rational):

so you want to solve \[\cos x -\sin x = 1\]

OpenStudy (rational):

the trick here is to use the identity \(\cos A\cos B-\sin A\sin B = \cos(A+B)\)

OpenStudy (rational):

familiar with that identity ?

OpenStudy (anonymous):

Yes, how would I apply it to cosx-sinx=1 though?

OpenStudy (rational):

Yes that's the interesting part, divide \(\sqrt{2}\) through out recalling the fact that \[\sin(45)=\cos(45)=\frac{1}{\sqrt{2}}\]

OpenStudy (rational):

\[\cos x -\sin x = 1\] dividing \(\sqrt{2}\) through out gives \[\cos x\frac{1}{\sqrt{2}} -\sin x\frac{1}{\sqrt{2}} = \frac{1}{\sqrt{2}}\] replace \(\frac{1}{\sqrt{2}}\) by \(\cos(45)\) or \(\sin(45)\) as needed : \[\cos x\cos(45) -\sin x\sin(45) = \frac{1}{\sqrt{2}}\]

OpenStudy (rational):

Now the left hand side is ready to use that identity, yes ?

OpenStudy (anonymous):

Got it. Why choose the square root of 2 specifically to divide though? I'm a little confused on that part.

OpenStudy (rational):

ikr, that \(\sqrt{2}\) might seem coming from nowhere, but there is a more general formula... maybe for now keep going and finish solving the equation

OpenStudy (anonymous):

Alright thank you so much for the help, I think I know how to solve it from here (:

OpenStudy (rational):

Nice :) about that \(\sqrt{2}\) : when you have an expression like \[a\cos x+b\sin x\] draw a right triangle with legs \(a,b\) 

OpenStudy (rational):

Notice that \(\color{blue}{\sin\theta=\frac{a}{\sqrt{a^2+b^2}}}\) and \(\color{purple}{\cos\theta=\frac{b}{\sqrt{a^2+b^2}}}\) , and the earlier expression can be rearranged as \[\begin{align}a\cos x+b\sin x&=\sqrt{a^2+b^2}\left(\color{blue}{\frac{a}{\sqrt{a^2+b^2}}}\cos x+\color{purple}{\frac{b}{\sqrt{a^2+b^2}}}\sin x\right)\\~\\ &=\sqrt{a^2+b^2}\left(\color{blue}{\sin\theta} \cos x+\color{purple}{\cos\theta}\sin x\right)\\~\\ &=\sqrt{a^2+b^2}\sin(\theta+x) \end{align}\]